# Domain Generalization in-the-Wild: Disentangling Classification from Domain-Aware Representations

### v1

@@ 5.1 Finetuning original pretrained CLIP
Figure 8: Comparison against more baselines
Figure 8 provides a broader comparison against additional baselines. We observe that conventional domain adversarial learning (DANN ganin2016domain_dann ), is harmful for CLIP, showing inferior performance compared to regular finetuning. This shows the potential disadvantage of enforcing domain invariance across the entire image encoder, which can lead to excessive forgetting of features learned during pretraining. This suggests the importance of approaches such as our proposed learning of targeted invariance through disentanglement.
Interestingly, on the most extremely OOD datasets, parameter-efficient finetuning (PEFT) techniques like CoOp zhou2022learning_coop and CLIP-Adapter gao2024clipadapter perform best. PEFT methods minimally change a small subset of the original CLIP weights. Consequently, their performance shows much lower variance across the datasets, with improvements (around 1-2%). It is important to note that on extreme OOD datasets, all end-to-end finetuning methods exhibit lower performance than the zero-shot CLIP baseline. While CLIP-DCA mitigates this performance drop compared to standard finetuning, it does not entirely overcome it.
This strong zero-shot performance has often been attributed to CLIP’s inherent OOD generalization capability. However, the study by mayilvahanan2024search challenges this assumption and shows that this generalization could be attributed to domain contamination. They show that when CLIP is retrained solely on natural images, its OOD performance drops to similar levels as models trained exclusively on ImageNet. This drop could offer a plausible explanation for observations like those motivating Wise-FT wortsman2022robust_wise , where standard finetuning was found to degrade OOD performance.
@@
Figure 9: Comparison against baselines after unlearning
5.2 Finetuning after unlearning
To further investigate the impact of potential domain contamination and to establish a more rigorously "unseen" evaluation, we applied the unlearning procedure detailed in Section 4.3 to the pretrained CLIP model. We then finetuned this "unlearned" model on ImageNet-1K and evaluated its performance. Table 2 shows the accuracies on the ImageNet variant datasets. For this analysis, we also include several end-to-end robust finetuning methods that add a linear classifier to CLIP. Due to their architecture, these specific baselines are evaluated only on the ImageNet variants as they cannot be adapted to datasets with different class labels.
Our results show that robust end-to-end finetuning methods remain effective for datasets that are less OOD even after unlearning. For instance, MIRO cha2022domain and Wise-FT wortsman2022robust_wise outperform regular finetuning on ImageNet-V1, ImageNet-V2, and ImageNet-Sketch. However, consistent with the trends seen with the non-unlearned model, performance significantly drops on datasets with larger OOD scores, such as ImageNet-A and ImageNet-R. Similarly, PEFT methods show slight improvements over the unlearned zero-shot baseline on ImageNet-V1, V2, and Sketch, but their performance drops on ImageNet-A and R.
@@ 5.3 Ablations
We study the contribution of the GCC data as shown in Table 3. A key observation is that the inclusion of GCC provides a notable benefit even for standard finetuning (FLYP) goyal2023finetune_flyp . This shows the general benefit of incorporating diverse, captioned data during finetuning. Given these benefits, an alternative or complementary approach could involve using MLLMs to generate rich textual descriptions for classes or images within the primary source dataset, similar to strategies explored in pratt2023does ; maniparambil2023enhancing , which use an LLM to describe class names. Despite the general improvements, our method consistently shows higher performance even when the GCC dataset was not included.
@@
Different components of CLIP-DCA. We study the effect of the different components of CLIP-DCA, as shown in Table 4. We isolate the use of domain descriptions from diffusion images to train the image domain head, the disentanglement loss between the class and domain heads to encourage invariance at the classifier, and the use of MLLM hidden states to encourage domain awareness in the text encoder. Simply introducing domain descriptions to make the image encoder aware of styles, without enforcing disentanglement at the classifier, shows only a marginal improvement over the FLYP baseline, suggesting that domain awareness alone is insufficient without a mechanism to disentangle classification from it, as CLIP may otherwise struggle to disregard domain-specific features irrelevant to classification. When we incorporate the disentanglement loss to encourage domain invariance at the decision-making layer, even without explicit domain awareness in the text encoder, performance slightly improves. This is further evidence for our core hypothesis that enabling the model to disregard domain-specific features during classification is important. Attempting to make both encoders domain-aware without the disentanglement loss results in no improvement over the baseline, indicating that awareness without a mechanism for invariance can be ineffective for OOD data.
@@
Limitations. One concern might be the reliance on synthetically generated diffusion images and MLLM-extracted features for domain awareness. However, this is mitigated by: (1) the small size of the diffusion dataset (4096 samples), (2) images synthesized using generic, class-agnostic style prompts, and (3) the MLLM processing multiple style-consistent images, which focuses it on style over objects. Furthermore, DANN ganin2016domain_dann and our ablations without disentanglement (Table 4), even with such data, fails to improve CLIP’s OOD performance (Table 3).
@@
The role of the MLLM may also be questionable, as LLaVA internally uses a CLIP-L encoder. However, LLaVA’s poor zero-shot image classification performance (Table 4), a known issue attributed to MLLMs’ improper alignment for classification zhang2024visually , justifies not using it as a direct classifier. Instead, we use an MLLM because CLIP captures global information from images, which prioritizes overall style tong2024eyes , making its representations suitable for domain-level information. The MLLM, with its language capabilities, is then able to explain the perceived domain styles into textual descriptions and provide informative hidden state representations.
@@
Lastly, our unlearning strategy involves making DomainNet images and random noise indistinguishable, differing from the standard approach sepahvandselective where samples are typically mapped to known OOD data. This adaptation was necessary as CLIP’s extensive web-scale pretraining makes finding truly unseen data challenging. Future work could explore more sophisticated unlearning methods for DG in-the-wild evaluation. Nevertheless, the significant degradation observed in zero-shot performance post-unlearning, and the fact that PEFT methods showed improvements on less OOD data but poorer performance on more OOD data, is evidence that our unlearning procedure functioned as intended.

### v2

@@ 4.1 Finetuning original pretrained CLIP
Figure 8: Comparison against more baselines.
Figure 8 provides a broader comparison against additional baselines. We observe that conventional domain adversarial learning (DANN (Ganin et al., 2016)), is harmful for CLIP, showing inferior performance compared to regular finetuning. This shows the potential disadvantage of enforcing domain invariance across the entire image encoder, which can lead to excessive forgetting of features learned during pretraining. This suggests the importance of approaches such as our proposed learning of targeted invariance through disentanglement.
Interestingly, on the most extremely OOD datasets, parameter-efficient finetuning (PEFT) techniques like CoOp (Zhou et al., 2022c) and CLIP-Adapter (Gao et al., 2024) perform best. PEFT methods minimally change a small subset of the original CLIP weights. Consequently, their performance shows much lower variance across the datasets, with improvements (around 1-2%). It is important to note that on extreme OOD datasets, all end-to-end finetuning methods exhibit lower performance than the zero-shot CLIP baseline. While CLIP-DCA mitigates this performance drop compared to standard finetuning, it does not entirely overcome it.
This strong zero-shot performance has often been attributed to CLIP’s inherent OOD generalization capability. However, the study by (Mayilvahanan et al., 2024) challenges this assumption and shows that this generalization could be attributed to domain contamination. They show that when CLIP is retrained solely on natural images, its OOD performance drops to similar levels as models trained exclusively on ImageNet. This drop could offer a plausible explanation for observations like those motivating Wise-FT (Wortsman et al., 2022), where standard finetuning was found to degrade OOD performance.
4.2 Finetuning after unlearning
To further investigate the impact of potential domain contamination and to establish a more rigorously "unseen" evaluation, we applied the unlearning procedure detailed in Section 3.3 to the pretrained CLIP model. We then finetuned this "unlearned" model on ImageNet-1K and evaluated its performance. Table 2 shows the accuracies on the ImageNet variant datasets. Full per-dataset accuracy details for all methods, both before and after unlearning, are provided in Appendix I. For this analysis, we also include several end-to-end robust finetuning methods that add a linear classifier to CLIP. Due to their architecture, these specific baselines are evaluated only on the ImageNet variants as they cannot be adapted to datasets with different class labels.
Our results show that robust end-to-end finetuning methods remain effective for datasets that are less OOD even after unlearning. For instance, MIRO (Cha et al., 2022) and Wise-FT (Wortsman et al., 2022) outperform regular finetuning on ImageNet-V1, ImageNet-V2, and ImageNet-Sketch. To broaden our comparison, we also include other recent CLIP-based DG methods such as MMA (Yang et al., 2024) and LwEIB (Yang et al., 2025), which similarly demonstrate that performance does not consistently generalize to more OOD datasets. However, consistent with the trends seen with the non-unlearned model, performance significantly drops on datasets with larger OOD scores, such as ImageNet-A and ImageNet-R. Similarly, PEFT methods show slight improvements over zero-shot on ImageNet-V1, V2, and Sketch, but their performance drops on ImageNet-A and R.
@@ 4.2 Finetuning after unlearning
Figure 9: Comparison against baselines after unlearning.
@@ 4.3 Ablations
We study the contribution of the GCC data as shown in Table 3. A key observation is that the inclusion of GCC provides a notable benefit even for standard finetuning (FLYP) (Goyal et al., 2023). This shows the general benefit of incorporating diverse, captioned data during finetuning. Given these benefits, an alternative or complementary approach could involve using MLLMs to generate rich textual descriptions for classes or images within the primary source dataset, similar to strategies explored in (Pratt et al., 2023; Maniparambil et al., 2023), which use an LLM to describe class names. Despite the general improvements, our method consistently shows higher performance even when the GCC dataset was not included.
@@
Different components of CLIP-DCA. We study the effect of the different components of CLIP-DCA, as shown in Table 4. We isolate the use of domain descriptions from diffusion images to train the image domain head, the disentanglement loss between the class and domain heads to encourage invariance at the classifier, and the use of MLLM hidden states to encourage domain awareness in the text encoder. Simply introducing domain descriptions to make the image encoder aware of styles, without enforcing disentanglement at the classifier, shows only a marginal improvement over the FLYP baseline, suggesting that domain awareness alone is insufficient without a mechanism to disentangle classification from it, as CLIP may otherwise struggle to disregard domain-specific features irrelevant to classification. When we incorporate the disentanglement loss to encourage domain invariance at the decision-making layer, even without explicit domain awareness in the text encoder, performance slightly improves. This is further evidence for our core hypothesis that enabling the model to disregard domain-specific features during classification is important. Attempting to make both encoders domain-aware without the disentanglement loss results in no improvement over the baseline, indicating that awareness without a mechanism for invariance can be ineffective for OOD data. To study the effect of the MLLM’s scale, we replaced LLaVA-8B with Gemini-2.5-Pro and observed a marginal performance difference (see Table˜H.1 in the Appendix), suggesting our method’s efficacy is not primarily dependent on the MLLM’s size but rather on the disentanglement framework itself. These results strongly support our central hypothesis: the significant performance gain of our full method demonstrates that the balance between domain awareness and disentangled invariance is the critical factor for robust generalization in this challenging setting.
@@
Limitations. One concern might be the reliance on synthetically generated diffusion images and MLLM-extracted features for domain awareness. However, this is mitigated by: (1) the small size of the diffusion dataset (4096 samples), (2) images synthesized using generic, class-agnostic style prompts, and (3) the MLLM processing multiple style-consistent images, which focuses it on style over objects. Furthermore, DANN (Ganin et al., 2016) and our ablations without disentanglement (Table 4), even with such data, fails to improve CLIP’s OOD performance (Table 3).
@@
The role of the MLLM may also be questionable, as LLaVA internally uses a CLIP-L encoder. However, LLaVA’s poor zero-shot image classification performance (Table 4), a known issue attributed to MLLMs’ improper alignment for classification (Zhang et al., 2024), justifies not using it as a direct classifier. Instead, we use an MLLM because CLIP captures global information from images, which prioritizes overall style (Tong et al., 2024), making its representations suitable for domain-level information. The MLLM, with its language capabilities, is then able to explain the perceived domain styles into textual descriptions and provide informative hidden state representations.
@@
Lastly, our unlearning strategy involves making DomainNet images and random noise indistinguishable, differing from Sepahvand et al. (2025) where samples are typically mapped to known OOD data. This adaptation was necessary as CLIP’s extensive web-scale pretraining makes finding truly unseen data challenging. Future work could explore more sophisticated unlearning methods for DG in-the-wild evaluation. Nevertheless, the significant degradation observed in zero-shot performance post-unlearning, and the fact that PEFT methods showed improvements on less OOD data but poorer performance on more OOD data, is evidence that our unlearning procedure functioned as intended.